Integrals Giving Inverse Trigonometric Functions


∫(xdx)/(25+16x⁴)


1
Expert's answer
2021-04-15T07:40:28-0400

"\displaystyle\int\frac{xdx}{25+16x^4}"


Note that if

"d(x^2)=(x^2)'dx=2xdx"

then

"xdx=\dfrac{1}{2}\cdot d(x^2)"


"\displaystyle\int\frac{xdx}{25+16x^4}="


"\displaystyle\int\frac{\frac{1}{2}\cdot d(x^2)}{25+16x^4}="


"\displaystyle\frac{1}{2}\cdot \int\frac{d(x^2)}{25+16x^4}="


"\displaystyle\frac{1}{2}\cdot \int\frac{d(x^2)}{25+16(x^2)^2}="


"\displaystyle\frac{1}{2}\cdot \int\frac{d(x^2)}{16\cdot(\frac{25}{16}+(x^2)^2)}="


"\displaystyle\frac{1}{2\cdot16}\cdot \int\frac{d(x^2)}{\frac{25}{16}+(x^2)^2}="


"\displaystyle\frac{1}{32}\cdot \int\frac{d(x^2)}{\frac{25}{16}+(x^2)^2}=" use antiderivative "\displaystyle\int\frac{dt}{a^2+t^2}=\frac{1}{a}\arctan\left(\frac{t}{a}\right)+c"


substituting "t=x^2" and "a^2=\dfrac{25}{16}"


"\displaystyle\frac{1}{32}\cdot\frac{1}{\sqrt{\frac{25}{16}}}\cdot\arctan\left(\frac{x^2}{\sqrt{\frac{25}{16}}}\right)+c="


"\displaystyle\frac{1}{32}\cdot\frac{1}{\frac{5}{4}}\cdot\arctan\left(\frac{x^2}{\frac{5}{4}}\right)+c="


"\displaystyle\frac{1}{32}\cdot\frac{4}{5}\cdot\arctan\left(\frac{4x^2}{5}\right)+c="


"\displaystyle\frac{1}{8}\cdot\frac{1}{5}\cdot\arctan\left(\frac{4x^2}{5}\right)+c="


"\displaystyle\frac{1}{40}\cdot\arctan\left(\frac{4x^2}{5}\right)+c"